OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? If the radius of a sphere is tripled, then its volume will increase 27 times. True False

OpenStudy (sleepyjess):

lets test it out say our radius of the first sphere is 5 second sphere is 15

OpenStudy (sleepyjess):

since we just did one with a radius of 5 we know that the volume of r=5 is 523.3333333.... now use 15 for r \(\dfrac{4}{3}\pi 15^3\)

OpenStudy (anonymous):

@sleepyjess i got 14130 if i was doing it right...

OpenStudy (sleepyjess):

\(15^3=3375\\3375*\pi=10597.5\\10597.5*(4/3)=14129.999999=14130\) so now do 523.3333*27 to see if it is equal 523.3333333*27=14129.9999999=14130

OpenStudy (anonymous):

@sleepyjess what would i do from there ? Is it true or false ?

OpenStudy (sleepyjess):

it is true since they are equal